OpenStudy (anonymous):

O(n) is an odd-number function.Find O(18), where O(n)=2n-1

OpenStudy (misty1212):

i have no idea what "odd number function" means but if \[O(n)=2n-1\] then \[O(18)=2\times 18-1\]

OpenStudy (anonymous):

Wait so what's the answer?

OpenStudy (anonymous):

Misty1212 wouldn't it be 35?

OpenStudy (mathmate):

Yes, 35 is correct! 2*18-1=36-1=35

OpenStudy (anonymous):

Thanks (:

OpenStudy (mathmate):

You're welcome! :)